OpenStudy (anonymous):

We started working on probability and I am having a hard time understanding how to do odds

OpenStudy (anonymous):

The odds in favor = success/ fail = 3 : 8

OpenStudy (anonymous):

I thought so, I have a few more problems do you think you could help me?

OpenStudy (anonymous):

Just post all up. There's so many talented helpers around you!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

If the probability an event will occur is 1/2 determine A) the probability that the event will fail B) the odds against the event occuring C) the odds in favor of the event occuring

OpenStudy (amistre64):

P(s) + P(f) = 1 .5 + P(f) = 1 ; when P(f) = .5

OpenStudy (amistre64):

in any given event; the probability associated with 1/2 is if there is a 1 to 1 ration of successes and failures; so the odds would be 1:1

OpenStudy (anonymous):

ok so is that the ans for all 3 problems or just A Thank you so much

OpenStudy (anonymous):

The key is the terms: P (Pass + Fail) = P(occur + not occur) = P(against + favor) = 1

OpenStudy (anonymous):

I understand the equation what I am having a hard time with is know where to place the values

OpenStudy (anonymous):

a card is chosen from a standard deck of cards determine the odds for and against picking a heart

OpenStudy (anonymous):

I know a deck of card is 52, but how many heart is it? P ( agaisnt heart) = number card of not heart/ number card of heart

OpenStudy (anonymous):

13 hearts total

OpenStudy (anonymous):

=> 52 - 13 = 39 cards not heart Thus P (against heart) = 39 : 13

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

a box contains 9 red and 2 blue marbles, if you selected one marble at random from the box determine the odds against selecting a red marble and explain how you determined your answer.

OpenStudy (anonymous):

P ( against red) = 2 : 9